# Prompting Techniques for Reducing Social Bias in LLMs through System 1 and System 2 Cognitive Processes

### v3

@@ 6 Does CoT Prompting Best Model System 2?
Now we further investigate whether CoT prompting is most similar to the way that LLMs model System 2 reasoning. While Figure 2 shows that the stereotyping rate of CoT is most similar to System 1 prompts, these may be from different test items. Here we tackle this question directly by computing the Kendall $\tau$ coefficient (Kendall, 1938) between CoT-prompted responses and those of the other variants. We use the Kendall $\tau$ ranked correlation because there is a natural order to anti-stereotypical, neutral, and stereotypical categorical values in our datasets. Table 3 lists these results. From this, we find that CoT prompting is most similar to the Standard zero-shot prompt, followed by System 1 prompting. In fact, it is most dissimilar to System 2 prompting. This pattern holds for the Human Persona and Machine Persona variants, where CoT is least correlated with the System 2 prompt variant.
@@
7 Invalid LLMs Responses
@@
8 Conclusion
Our study has contributed to the understanding and reduction of social biases in LLMs through prompting techniques grounded in dual process theory. By harnessing the cognitive frameworks of System 1 and System 2, as well as the incorporation of human-like personas and debiasing prompts, our research not only clarifies the role of these cognitive processes in LLMs but also demonstrates practical methods for reducing biases. Our findings reveal that System 2 prompts, particularly when combined with a Human Persona, consistently reduce stereotypical judgments across various social bias categories. Biases were further reduced when using a debiasing prompt, which can be seen as a social bias-focused System 2 prompt, along with a human persona. This indicates a profound potential for combining analytical thought processes and personalized prompting to enhance the ethical performance of LLMs. Furthermore, our use of different models and bias datasets has allowed us to explore the diverse applications of these techniques, ensuring our results are robust and applicable across different contexts. These findings underscore the potential of sophisticated prompting strategies in enhancing the ethical aspects of AI, pointing towards a future where LLMs can assist in creating more inclusive digital environments. Through continued exploration and refinement of these methods, we anticipate further advancements in the responsible deployment of AI technologies.

### v4

@@ Appendix D Does CoT Prompting Best Model System 2?
Now we further investigate whether CoT prompting is most similar to the way that LLMs model System 2 reasoning. While Figure 2 shows that the stereotyping rate of CoT is most similar to System 1 and Standard prompts, these may be from different test items. Here we tackle this question directly by computing the Kendall $\tau$ coefficient (Kendall, 1938) between CoT-prompted responses and those of the other variants. We use the Kendall $\tau$ ranked correlation because there is a natural order to anti-stereotypical, neutral, and stereotypical categorical values in our datasets. Table 10 lists these results. From this, we find that CoT prompting is most similar to the Standard zero-shot prompt, followed by System 1 prompting. In fact, it is most dissimilar to System 2 prompting. This pattern holds for the Human Persona and Machine Persona variants, where CoT is least correlated with the System 2 prompt variant.
@@
Appendix E Invalid LLMs Responses